OpenStudy (anonymous):

How many solutions are there to the following system of equations? 4x – 14y = 6 –2x + 7y = –3 0 1 2 infinitely many

OpenStudy (campbell_st):

well take the 2nd equation and multiply it by -2... what do you get..?

OpenStudy (anonymous):

How would I multiply it exactly? o.o

OpenStudy (campbell_st):

well just double each term and change the sign to the opposite of whats shown

OpenStudy (anonymous):

still confuse

OpenStudy (anonymous):

-2x*-2=4x 7y*-2=-14y -3*-2=6 So, the equations are the same. Now, use logic to get your answer.

OpenStudy (anonymous):

okay, that makes sense, but can you give me one more step? I'm not quite sure where to take this @curiouscrab

OpenStudy (jdoe0001):

hmm @Levi565 can you solve each for "y"?

OpenStudy (jdoe0001):

what would you get?

OpenStudy (anonymous):

How many numbers can you replace x or y with and still have both equations equal? Well since both equations are already equal, then...

OpenStudy (anonymous):

Uhhhm.........no they wouldnt be equal

OpenStudy (anonymous):

Yes they are. It's very hard to explain.

OpenStudy (ranga):

These two equations are essentially the same. If you multiply the 2nd equation by -2 you get the first equation. So we have two unknowns: x and y. But only one equation. That means you can choose any value for one unknown and use the one equation to solve for the other unknown. And you can keep doing this forever. Therefore, this system has ??? solutions.

OpenStudy (anonymous):

4x - 14y = 6 (change to slope intercept form ) -14y = -4x + 6 y = -4x/-14 + 6/-14 y = 2/7x - 3/7 -2x + 7y = -3 (change to slope intercept form) 7y = 2x - 3 y = 2/7x - 3/7 since both equations are the same in slope intercept form, then there are INFINITE SOLUTIONS